how much 12% acid solution and pure water must be mixed to produce 19.2 gallons of a mixture that is 7% acid?
Answer by lwsshak3(11628)

how much 12% acid solution and pure water must be mixed to produce 19.2 gallons of a mixture that is 7% acid?
**
let x= gallons of pure water to be mixed
19.2-x= gallons of 12% solution to be mixed
..
12%(19.2-x)+0%x=7%*19.2
2.304-.12x+0=1.344
.12x=.96
x=8
gallons of pure water to be mixed=8
gallons of 12% solution to be mixed=11.2